# Uncertainty-Aware Diffusion Guided Refinement of 3D Scenes

### v1

@@ Abstract
Reconstructing 3D scenes from a single image is a fundamentally ill-posed task due to the severely under-constrained nature of the problem. Consequently, when the scene is rendered from novel camera views, existing single image to 3D reconstruction methods render incoherent and blurry views. This problem is exacerbated when the unseen regions are far away from the input camera. In this work, we address these inherent limitations in existing single image-to-3D scene feedforward networks. To alleviate the poor performance due to insufficient information beyond the input image’s view, we leverage a strong generative prior in the form of a pre-trained latent video diffusion model, for iterative refinement of a coarse scene represented by optimizable Gaussian parameters. To ensure that the style and texture of the generated images align with that of the input image, we incorporate on-the-fly Fourier-style transfer between the generated images and the input image. Additionally, we design a semantic uncertainty quantification module that calculates the per-pixel entropy and yields uncertainty maps used to guide the refinement process from the most confident pixels while discarding the remaining highly uncertain ones. We conduct extensive experiments on real-world scene datasets, including in-domain RealEstate-10K and out-of-domain KITTI-v2, showing that our approach can provide more realistic and high-fidelity novel view synthesis results compared to existing state-of-the-art methods.
@@ 1 Introduction
One way to compensate for this lack of multi-view information is to use generative priors so that they can provide valuable complementary information. The principle is to synthesize plausible 3D details even for regions not visible in the image, where plausibility refers to the fact that visible portions should remain multi-view consistent, while occluded or out-of-view areas should integrate seamlessly with the context of the scene (refer to supplementary for qualitative results). Some existing approaches toward solving this problem include text-guided inpainting [13] and distillation of 2D diffusion priors [14]. However, these methods either lack 3D correspondence awareness [15] or do not have any temporal continuity as they denoise each image separately. This leads to issues with multi‐view consistency for large viewpoint changes as the camera moves further away from the source. One way to address this challenge is to reformulate the problem as a 2.5D reconstruction [16, 17], which can be limited in terms of fully capturing unseen regions. Moreover, extending such approaches to fully unbounded real world scenes, where the scene may not loop back to the starting point, is non-trivial. These methods often rely heavily on generated features, leading to mode collapse [18], content distortion [18, 19] and visible artifacts in rendered views [20, 21], especially when the camera motion spans a large range. Although diffusion priors are well-suited for object‐level enhancement, these limitations indicate that priors designed for object reconstruction, cannot be directly applied for scene refinement.
@@
To this end, we propose Uncertainty-Aware Diffusion-Guided Refinement of 3D Scenes: UAR-Scenes, an uncertainty‐aware Gaussian splatting based refinement framework that combines the strengths of regressive and generative approaches. We perform the refinement using Adaptive Densification and Pruning (ADP) [6], a mechanism that controls the density of 3D points by introducing or deleting Gaussian primitives (check supplementary Section 6.1). ADP plays a crucial role in expanding or shrinking Gaussians depending on whether the region is over-constructed or not. Therefore, in observed regions where the confidence per pixel is high, our method mostly preserves Gaussian primitives with high‐fidelity details obtained from the reconstruction backbone, while in unobserved or occluded areas, it learns to introduce new Gaussians by using uncertainty-driven sampling from a Latent Video Diffusion Model (LVDM) [22] to produce plausible scene completions. We leverage an MLLM [23] along with an open-vocabulary segmentation model [24] to compute the per-pixel entropy present in LVDM generated views which gives us the per-pixel uncertainty associated with the generated images. In this way, the inherent ambiguity present in single-view reconstruction methods can be addressed, expanding the applicability of our pipeline from simple object‐centric datasets or scenes with limited backgrounds [25] to arbitrarily large real‐world unbounded scenes [26]. Furthermore, it does not rely on any form of ground truth supervision (2D or 3D), making the entire process self‐supervised and thus, post-hoc compatible with any existing 3D reconstruction algorithm. This design effectively addresses ambiguities, such as missing artifacts and inconsistent 3D reconstructions inherent in single‐view settings. Thus, in a nutshell, our proposed model integrates a differentiable feed-forward scene reconstruction pipeline that operates from a single image with an uncertainty-aware LVDM for self-supervised refinement of noisy scenes.
@@
  • To address any texture and style inconsistencies between the input image and the generated pseudo views from the LVDM, we design a Fourier Style Transfer (FST)-based alignment methodology.
@@ 2 Related Works
Regressive 3D Reconstruction from Sparse Views. Existing generalizable reconstruction methods [1, 28, 12, 29, 30, 31] aim to train a single feed-forward model which can predict a 3D scene directly. Such methods include PixelSplat [11] which utilizes epipolar transformers to reduce scale ambiguity and obtain better scene features. Similarly, MVSplat [12] builds cost volumes to obtain Gaussians in a fast and efficient manner. LatentSplat [30] utilizes a GAN decoder as a generative component but GANs are known to be highly unstable [32] and hence unsuitable for modeling real-world unbounded scenes. Recently, Flash3D [1] introduced a single image to 3D scene pipeline leveraging a pre-trained depth estimator. In this paper, we build upon this regressive framework and show how most works falling in this paradigm fail to generate realistic completions for unseen regions.
@@
Generative Scene Reconstruction and Synthesis. Our method falls within an under-explored area of utilizing only 2D (pseudo) labels to supervise 3D tasks using latent diffusion models. Existing approaches like [33, 34, 35] denoise multi-view images using a 3D-aware denoiser. However, such methods do not produce a coherent 3D structure since the denoising process is done separately for each image leading to sub-optimal multi-view consistency. Works such as 3DGS-Enhancer [36] address a similar problem in a sparse view setting but they employ a mix of refined views along with un-refined views when training the video-diffusion model. Other works such as CAT3D [37], GeNVS [38], ReconFusion [14], LM-Gaussian [39], and ReconX [40] [^1] Code and models are not publicly available for these methods. ReconFusion was trained on an internal dataset. all tackle similar problems but they deal with the problem on an object level by employing a 2D diffusion model [41] which is not multi-view consistent. Further, none of these methods optimize the Gaussians directly but instead fine-tune an existing generative diffusion model on a dataset to make it generalizable, which is expensive, and ties the models down to the particular distribution on which it is trained.
@@
Uncertainity Quantification in 3D Reconstruction. Quantifying uncertainty has received relatively less attention in 3D scene reconstruction, especially considering it’s importance in the context of single image to 3D reconstruction where there is inherently a lot of ambiguity. One of the first approaches in estimating epistemic uncertainty in NeRFs from sparse views using variational inference was studied in works [42, 43, 44] but owing to increasing parameters alongwith using different models, these approaches are inherently expensive. BayesRays [45] proposed a method to quantify uncertainty in any NeRF based reconstruction module in a post-hoc fashion. Further, uncertainty has also been explored in the context of active-view selection and path planning in the field of 3D mapping in robotics [46, 7] to estimate which next view maximizes the information gain over a given pre-trained model by approximating the hessian matrix of the log-likelihood function. However, this becomes un-scalable to compute for each Gaussian primitive. Inspired by these approaches, we opt for a semantic scene understanding based approach for quantifying the uncertainty in novel test-time views for refinement.
@@
Figure 2: Workflow of UAR-Scenes. For a conditioning image $\mathcal{I}$, a pre-trained 3D reconstruction model $\mathcal{F(\cdot)}$ produces coarse gaussians ($\gamma_{n}$) representing the scene $\phi$, represented by optimizable gaussian parameters. Using temporally consistent pseudo 2D supervisory images (~$\widetilde{\mathcal{I}}_{\text{p}}$) sampled from the pre-trained camera extrinsic embedded LVDM model [22], we iteratively refine the gaussians $\gamma_{n}$ using Adaptive Densification and Pruning (ADP) [47] to obtain clean gaussians ($\gamma^{{}^{\prime}}_{n}$). To gauge the uncertainty of each pixel p in the generated pseudo images (~$\widetilde{\mathcal{I}}_{\text{p}}$), we propose a semantic uncertainty quantification method. We estimate the entropy present in each (~$\widetilde{\mathcal{I}}_{\text{p}}$) obtained by utilizing an off-the-shelf open-vocabulary segmentation model $\mathcal{S}$ [24] using which we obtain uncertainty maps $\mathcal{U}$ (as shown in Figure 3). We take the hadamard product between ~$\widetilde{\mathcal{I}}_{\text{p}}$ and $\mathcal{U}$ forming the target objective for the refinement loss in Equation 6, which guides the refinement process.
@@ 3 Methodology
Overview. We provide an overview of our uncertainty-sensitive refinement pipeline in Figure 2. To get an accurate target representation of the scene, the geometric features from the backbone deterministic reconstruction model are preserved and the textures from the LVDM are utilized. The discussion starts with the deterministic model, serving as the initial point for refining scene Gaussians, detailed in Section 3.1. Then, we briefly discuss the topic of camera controllable LVDMs in Section 3.2 since it is crucial to align the pseudo views and the noisy rendered views so that they are in the same camera space. The next part addresses the method of weighting the optimization objective based on the confidence associated with each pixel, as outlined in Section 3.3. The final section describes the application of a video diffusion model to generate pseudo-views and facilitate iterative refinement, further elaborated in Section 3.4.
@@ 3.1 Initial Coarse 3D Representation
[…]
@@ 3.2 LVDM with Camera Control
where $\mathcal{C}_{(R,T)}$ is the embedding of the camera extrinsics. Upon completion of the training process, users provide the conditioning image $y$ along with the camera’s extrinsics ($\mathbf{R},\mathbf{T}$), to generate the target views corresponding to the specified poses. For the sake of brevity, we refer to this version of camera controllable LVDM as LVDM ($\Phi$) unless mentioned otherwise.
@@ 3.3 Semantic Uncertainty Quantification
Figure 3: Uncertainty Map Estimation. On the right is the obtained uncertainty map from Pseudo Ground truth Image ~$\widetilde{\mathcal{I}}_{\text{p}}$ (on the left). It is generated by the LVDM after applying FST [52]. These maps are crucial for guiding the gaussian refinement process to progressively focus more on the confident (blue and green regions) and not on blurry ceiling and stairs (red).
@@
Uncertainty Distillation. For each pseudo-view ~$\widetilde{\mathcal{I}}_{\text{p}}$ generated by the LVDM, we use a Multimodal Large Language Model (MLLM) [23] to detect objects (even if partially visible) and obtain their object tags $O_{1\dots N}$. These $N{+}1$ classes (including background) are then passed to an open-vocabulary segmentation model $\mathcal{S}$ [24], producing pixel-level logits:
$o_{t}=\mathcal{S}\Bigl{(}\widetilde{\mathcal{I}}_{\text{p}},\,O\Bigr{)}\quad\in\mathbb{R}^{n\times(N+1)},$ (3)
@@
which forms the uncertainty map $\mathcal{U}\in\mathbb{R}^{H\times W}$ which is then subsequently used to perform pixel-wise multiplication with the LVDM image during refinement. Intuitively, $\mathcal{U}_{(i,j)}$ captures the uncertainty present in the segmentation map, whereby a higher entropy indicates a more uniform (i.e. less confident) class distribution for a pixel, while a lower entropy implies higher confidence in the predicted semantic classes. This entropy-based uncertainty measure thus provides a principled way to assess the reliability of the pseudo-view segmentation for our uncertainty-aware refinement. Further implementation details on how the MLLM was used for object tagging is explained in Section 6.4 of the supplementary material.
[…]
@@ 3.4 Iterative Refinement of Gaussians
After obtaining the coarse Gaussians $\gamma^{{}^{\prime}}_{n}$ from $\mathcal{F(\cdot)}$ in Section 3.1, utilizing camera controlled LVDM $\mathcal{G}$ for generating pseudo views for guidance (in Section 3.2) and defining our uncertainty weighted refinement objective in Section 3.3, we optimize the parameters of $\gamma_{n}$ iteratively to obtain $\gamma^{{}^{\prime}}_{n}$. However, it is observed that naive refinement of all the parameters leads to distorted renderings. We discuss the optimization strategy, style, texture alignment and the final refining objective for our pipeline below.
@@
Texture alignment with FST ($\Phi$). LVDMs tend to exhibit over-saturated colors and features which may become more pronounced in indoor scenes (as shown in Figure 5). While the generated images may be plausible, they might hurt the performance quantitatively in metrics like PSNR and SSIM which evaluate pixel-level accuracy. To address this, we perform Fourier Style Transfer (FST) adaptation [52] on the fly between the input image $\mathcal{I}$ and the output images generated by the LVDM model. This helps align the pseudo-images ~$\widetilde{\mathcal{I}}_{\text{p}}$ with the texture of the incoming source images. Aligning textures in this manner contributes to improving the overall performance, as shown in Table 4.

### v2

# Uncertainty-Aware Diffusion-Guided Refinement of 3D Scenes

@@ Abstract
Reconstructing 3D scenes from a single image is a fundamentally ill-posed task due to the severely under-constrained nature of the problem. Consequently, when the scene is rendered from novel camera views, existing single image to 3D reconstruction methods render incoherent and blurry views. This problem is exacerbated when the unseen regions are far away from the input camera. In this work, we address these inherent limitations in existing single image-to-3D scene feedforward networks. To alleviate the poor performance due to insufficient information beyond the input image’s view, we leverage a strong generative prior, in the form of a pre-trained latent video diffusion model, for iterative refinement of a coarse scene represented by optimizable Gaussian parameters. To ensure that the style and texture of the generated images align with that of the input image, we incorporate on-the-fly Fourier-style transfer between the generated images and the input image. Additionally, we design a semantic uncertainty quantification module that calculates the per-pixel entropy and yields uncertainty maps used to guide the refinement process from the most confident pixels while discarding the remaining highly uncertain ones. We conduct extensive experiments on real-world scene datasets, including in-domain RealEstate-10K and out-of-domain KITTI-v2, showing that our approach can provide more realistic and high-fidelity novel view synthesis results compared to existing state-of-the-art methods. The project page is available at: https://github.com/UCR-Vision-and-Learning-Group/UAR-Scenes
@@ 1 Introduction
One way to compensate for this lack of multi-view information is to use generative priors so that they can provide valuable complementary information. The principle is to synthesize plausible 3D details even for regions not visible in the image, where plausibility refers to the fact that visible portions should remain multi-view consistent, while occluded or out-of-view areas should integrate seamlessly with the context of the scene (refer to supplementary for qualitative results). Some existing approaches toward solving this problem include text-guided inpainting [13] and distillation of 2D diffusion priors [14]. However, these methods either lack 3D correspondence awareness [15] or do not have any temporal continuity as they denoise each image separately. This leads to issues with multi‐view consistency for large viewpoint changes as the camera moves further away from the source. One way to address this challenge is to reformulate the problem in a 2.5D reconstruction [16, 17] setting, which can be limited in terms of fully capturing unseen regions. Moreover, extending such approaches to fully unbounded real world scenes, where the scene may not loop back to the starting point, is non-trivial. These methods often rely heavily on just generated features, leading to mode collapse [18], content distortion [18, 19] and visible artifacts in rendered views [20, 21], especially when the camera motion spans a large range. Although diffusion priors are well-suited for object‐level enhancement, these limitations indicate that priors primarily designed for object reconstruction, cannot be directly applied for scene refinement.
@@
To this end, we propose Uncertainty-Aware Diffusion-Guided Refinement of 3D Scenes: UAR-Scenes, an uncertainty‐aware Gaussian splatting based refinement framework that combines the strengths of regressive and generative approaches. We perform the refinement using Adaptive Densification and Pruning (ADP) [6], a mechanism that controls the density of 3D points by introducing or deleting Gaussian primitives (check supplementary Section 6.1). ADP plays a crucial role in expanding or shrinking Gaussians depending on whether the region is over-constructed or not. Therefore, in observed regions where the confidence per pixel is high, our method mostly preserves Gaussian primitives with high‐fidelity details obtained from the reconstruction backbone, while in unobserved or occluded areas, it learns to introduce new Gaussians by using uncertainty-driven sampling from a Latent Video Diffusion Model (LVDM) [22] to produce plausible scene completions. We leverage an MLLM [23], along with an open-vocabulary segmentation model [24], to compute the per-pixel entropy present in LVDM generated views which gives us the per-pixel uncertainty associated with the generated images. In this way, the inherent ambiguity present in single-view reconstruction methods can be addressed, expanding the applicability of our pipeline from simple object‐centric datasets or scenes with limited backgrounds [25] to arbitrarily large real‐world unbounded scenes [26]. Furthermore, it does not rely on any form of ground truth supervision (2D or 3D), making the entire process self‐supervised and thus, post-hoc compatible with any existing 3D reconstruction algorithm. This design effectively addresses ambiguities, such as missing artifacts and inconsistent 3D reconstructions inherent in single‐view settings. Thus, in a nutshell, our proposed model integrates a differentiable feed-forward scene reconstruction pipeline that operates from a single image with an uncertainty-aware LVDM for self-supervised refinement of noisy scenes.
@@
  • To address any texture and style inconsistencies between the input image and the generated pseudo views from the LVDM, we propose a Fourier Style Transfer (FST)-based texture alignment methodology.
@@ 2 Related Works
Regressive 3D Reconstruction from Sparse Views. Existing generalizable reconstruction methods [1, 28, 12, 29, 30, 31, 32] aim to train a single feed-forward model which can predict a 3D scene directly. Such methods include PixelSplat [11] which utilizes epipolar transformers to reduce scale ambiguity and obtain better scene features. Similarly, MVSplat [12] builds cost volumes to obtain Gaussians in a fast and efficient manner. LatentSplat [30] utilizes a GAN decoder as a generative component but GANs are known to be highly unstable [33] and hence unsuitable for modeling real-world unbounded scenes. Recently, Flash3D [1] introduced a single image to 3D scene pipeline leveraging a pre-trained depth estimator. In this paper, we build upon this regressive framework and show how most works falling in this paradigm fail to generate realistic completions for unseen regions.
@@
Generative Scene Reconstruction and Synthesis. Our method falls within an under-explored area of utilizing only 2D (pseudo) labels to supervise 3D tasks using latent diffusion models. Existing approaches like [34, 35, 36, 37] denoise multi-view images using a 3D-aware denoiser. However, such methods do not produce a coherent 3D structure since the denoising process is done separately for each image leading to sub-optimal multi-view consistency. Works such as 3DGS-Enhancer [38] address a similar problem in a sparse view setting but they employ a mix of refined views along with un-refined views when training the video-diffusion model. Other works such as CAT3D [39], GeNVS [40], ReconFusion [14], LM-Gaussian [41], and ReconX [42] [^1] Code and models are not publicly available for these methods. ReconFusion was trained on an internal dataset. all tackle similar problems but they deal with the problem by employing expensive conditioning techniques [43] which may not always be multi-view consistent. Further, none of these methods optimize the Gaussians directly but instead fine-tune an existing generative diffusion model on datasets to make it generalizable, which ties the models down to the particular distribution on which it is trained.
@@
Uncertainty Quantification in 3D Reconstruction. Quantifying uncertainty has received relatively less attention in 3D scene reconstruction, especially considering it’s importance in the context of single image to 3D reconstruction where there is inherently a lot of ambiguity. One of the first approaches in estimating epistemic uncertainty in NeRFs from sparse views using variational inference was studied in works [44, 45, 46] but owing to increasing parameters alongwith using different models, these approaches are inherently expensive. BayesRays [47] proposed a method to quantify uncertainty in any NeRF based reconstruction module in a post-hoc fashion. Further, uncertainty has also been explored in the context of active-view selection and path planning in the field of 3D mapping in robotics [48, 7] to estimate which next view maximizes the information gain over a given pre-trained model by approximating the hessian matrix of the log-likelihood function. However, this becomes un-scalable to compute for each Gaussian primitive. Inspired by these approaches, we opt for a semantic scene understanding based approach for quantifying the uncertainty in novel test-time views for refinement.
@@
Figure 2: Workflow of UAR-Scenes. For a conditioning image $\mathcal{I}$, a pre-trained 3D reconstruction model $\mathcal{F(\cdot)}$ produces coarse gaussians ($\gamma_{n}$) representing the scene $\phi$, represented by optimizable gaussian parameters. Using temporally consistent pseudo 2D supervisory images ($\widetilde{\mathcal{I}}_{\text{p}}$) sampled from the pre-trained camera extrinsic embedded LVDM model [22], we iteratively refine the gaussians $\gamma_{n}$ using Adaptive Densification and Pruning (ADP) [49] to obtain clean gaussians ($\gamma^{{}^{\prime}}_{n}$). To gauge the uncertainty of each pixel p in the generated pseudo images ($\widetilde{\mathcal{I}}_{\text{p}}$), we propose a semantic uncertainty quantification method. We estimate the entropy present in each ($\widetilde{\mathcal{I}}_{\text{p}}$) obtained by utilizing an off-the-shelf open-vocabulary segmentation model $\mathcal{S}$ [24] using which we obtain uncertainty maps $\mathcal{U}$ (as shown in Figure 3). We take the hadamard product between $\widetilde{\mathcal{I}}_{\text{p}}$ and $\mathcal{U}$ forming the target objective for the refinement loss in Equation 6, which guides the refinement process.
@@ 3 Methodology
Overview. We provide an overview of our uncertainty-sensitive refinement pipeline in Figure 2. To get an accurate target representation of the scene, the geometric features from the backbone deterministic reconstruction model are preserved and the textures from the LVDM are utilized. The discussion starts with the deterministic model, serving as the initial point for refining scene Gaussians, detailed in Section 3.1. Then, we briefly discuss the topic of camera controllable LVDMs in Section 3.2 since it is crucial to align the pseudo views and the noisy rendered views so that they are in the same camera space. The next part addresses the method of weighting the optimization objective based on the confidence associated with each pixel, as outlined in Section 3.3. The final section describes the application of the video diffusion model to generate pseudo-views and facilitate iterative refinement, further elaborated in Section 3.4.
@@ 3.1 Initial Coarse 3D Representation
where $\mathcal$$\tilde${I}} is the image rendered by the initialized model $\mathcal{F(\cdot)}$ rasterizing its layered Gaussians and $\mathcal{I}_{\text{gt}}$ is the corresponding ground-truth image at train time. $\alpha$ is a hyper-parameter weighting the SSIM [53] loss.
@@ 3.2 LVDM with Camera Control
where $\mathcal{C}_{(R,T)}$ is the embedding of the camera extrinsics. Upon completion of the training process, the conditioning image $y$ can be provided along with the camera’s extrinsics ($\mathbf{R},\mathbf{T}$), to generate the target views corresponding to the specified poses. For the sake of brevity, we refer to this version of camera controllable LVDM as LVDM ($\mathcal{G}$) unless mentioned otherwise.
@@ 3.3 Semantic Uncertainty Quantification
Figure 3: Uncertainty Map Estimation. On the right is the obtained uncertainty map from Pseudo Ground truth Image $\widetilde{\mathcal{I}}_{\text{p}}$ (on the left). It is generated by the LVDM after applying FST [54]. These maps are crucial for guiding the Gaussian refinement process to progressively focus more on the confident (blue and green regions) and not on blurry ceiling and stairs (red).
@@
Uncertainty Distillation. For each pseudo-view $\widetilde{\mathcal{I}}_{\text{p}}$ generated by the LVDM, we use a Multimodal Large Language Model (MLLM) [23] to detect objects (even if partially visible) and obtain their object tags $O_{1\dots N}$. These N+1N{+}1 classes (including background) are then passed to an open-vocabulary segmentation model $\mathcal{S}$ [24], producing pixel-level logits:
$o_{t}=\mathcal{S}\Bigl(\widetilde{\mathcal{I}}_{\text{p}},\,O\Bigr)\in\mathbb{R}^{n\times(N+1)},$ (3)
@@
which forms the uncertainty map $\mathcal{U}\in\mathbb{R}^{H\times W}$ which is then subsequently used to perform pixel-wise multiplication with the LVDM image during refinement. Intuitively, $\mathcal{U}_{(i,j)}$ captures the uncertainty present in the segmentation map, whereby a higher entropy indicates a more uniform (i.e. less confident) class distribution for a pixel, while a lower entropy implies higher confidence in the predicted semantic classes. This entropy-based uncertainty measure thus provides a principled way to assess the reliability of the pseudo-view segmentation for our uncertainty-aware refinement. Further implementation details on how the MLLM was used for object tagging alongwith an open-set segmentation model [24] is explained in the supplementary.
Uncertainty-Weighted Reconstruction. Instead of relying solely on the standard reconstruction loss introduced in Section 3.1, we leverage the derived per-pixel uncertainty map $\mathcal{U}_{(i,j)}$ to guide the refinement process. In particular, given the rendered image $\mathcal$$\tilde${I}} from the pre-trained reconstruction model $\mathcal{F}(\cdot)$ and the pseudo image $\widetilde{\mathcal{I}}_{\text{p}}$, we modify Equation 1 to compute an uncertainty-weighted reconstruction loss by taking the Hadamard (pixel-wise) product between the reconstruction error and the complement of the uncertainty map to get the pixel-wise confidence:
@@ 3.4 Iterative Refinement of Gaussians
After obtaining the coarse Gaussians $\gamma_{n}$ from $\mathcal{F(\cdot)}$ in Section 3.1, utilizing camera controlled LVDM $\mathcal{G}$ for generating pseudo views for guidance (in Section 3.2) and defining our uncertainty weighted refinement objective in Section 3.3, we optimize the parameters of $\gamma_{n}$ iteratively to obtain $\gamma^{{}^{\prime}}_{n}$. However, it is observed that naive refinement of all the parameters leads to distorted renderings. We discuss the optimization strategy, style, texture alignment and the final refining objective for our pipeline below.
@@
Texture alignment with FST ($\Phi$). LVDMs tend to exhibit over-saturated colors and features which may become more pronounced in indoor scenes (as shown in Figure 5). While the generated images may be plausible, they might hurt the performance quantitatively in metrics like PSNR and SSIM which evaluate pixel-level accuracy. To address this, we perform Fourier Style Transfer (FST) adaptation [54] on the fly between the input image $\mathcal{I}$ and the output images generated by the LVDM model. This helps align the pseudo-images $\widetilde{\mathcal{I}}_{\text{p}}$ with the texture of the incoming source images. Aligning textures in this manner contributes to improving the overall performance, as shown in Table 4.